OpenStudy (anonymous):

Solve the system of equations. 5x - y = 3 and -10x + 2y = -6 A. (0, -3) B. (-3, 0) C. (3, 12) D. no solution E. infinitely many solutions

hero (hero):

5x - y = 3 -10x + 2y = -6 y = 5x - 3 y = (-6 + 10x)/2 y = y 2(5x -3) = 10x - 6 10x - 6 = 10x - 6 Infinite Solution